# Solving Football by Exploiting Equilibrium Structure of 2p0s Differential Games with One-Sided Information

### v3

@@ Abstract
For a two-player imperfect-information extensive-form game (IIEFG) with $K$ time steps and a player action space of size $U$, the game tree complexity is $U^{2K}$, causing existing IIEFG solvers to struggle with large or infinite $(U,K)$, e.g., differential games with continuous action spaces. To partially address this scalability challenge, we focus on an important class of 2p0s games where the informed player (P1) knows the payoff while the uninformed player (P2) only has a belief over the set of $I$ possible payoffs. Such games encompass a wide range of scenarios in sports, defense, cybersecurity, and finance. We prove that under mild conditions, P1’s (resp. P2’s) equilibrium strategy at any infostate concentrates on at most $I$ (resp. $I+1$) action prototypes. When $I\ll U$, this equilibrium structure causes the game tree complexity to collapse to $I^{K}$ for P1 when P2 plays pure best responses, and $(I+1)^{K}$ for P2 in a dual game where P1 plays pure best responses. We then show that exploiting this structure in standard learning modes, i.e., model-free multiagent reinforcement learning and model predictive control, is straightforward, leading to significant improvements in learning accuracy and efficiency from SOTA IIEFG solvers. Our demonstration solves a 22-player football game ($K=10$, $U=\infty$) where the attacking team has to strategically conceal their intention until a critical moment in order to exploit information advantage. Code is available here.
@@ 1 Introduction
The strength of game solvers has grown rapidly in the last decade, beating elite-level human players in Chess (Silver et al., 2017a), Go (Silver et al., 2017b), Poker (Brown & Sandholm, 2019; Brown et al., 2020b), Diplomacy (FAIR† et al., 2022), Stratego (Perolat et al., 2022), among others with increasing complexity. However, most of the existing solvers with proven convergence, either based on regret matching (Tammelin, 2014; Burch et al., 2014; Moravčík et al., 2017; Brown et al., 2020b; Lanctot et al., 2009) or gradient descent-ascent (McMahan, 2011; Perolat et al., 2021; Sokota et al., 2022; Cen et al., 2021; Vieillard et al., 2020), have computational complexities increasing with the size of the finite action set, and suffer from game-tree complexities growing exponentially with both the action size $U$ and the tree depth $K$. Real-world games, however, often have continuous actions and happen in continuous time and state spaces, making them differential in nature. Directly applying existing solvers to differential games would require either insightful action-state-time abstraction or enormous compute. Neither are readily available.
@@
In this paper, we address this scalability challenge for an important subset of 2p0s differential games where the informed player (P1) knows the payoff type while the uniformed player (P2) only holds a public belief $p_{0}\in\Delta(I)$ over the finite set of $I$ possible types. At the beginning of the game, nature draws a game type according to $p_{0}$ and informs P1 about the type. As the game progresses, the public belief $p$ about the true game type is updated from $p_{0}$ based on the action sequence taken by P1 and his strategy via the Bayes’ rule. P1’s (resp. P2’s) goal is to minimize (resp. maximize) the expected payoff over $p_{0}$. Due to the zero-sum nature, P1 may need to delay information release or manipulate P2’s belief to take full advantage of information asymmetry. While restricted, such games represent a wide range of attack-defense scenarios including football set-pieces where the attacker has private information about which play is to be executed, and missile defense where multiple potential targets are concerned. The setting of one-sided information, i.e., P1 knows everything about P2, is necessary for P2 to derive defense strategies in risk-sensitive games. We call this focused set of games “2p0s1”.
@@
  • We demonstrate how this structural knowledge can significantly accelerate game solving: (1) For value and policy approximation settings where the ground-truth NE is available, we achieve qualitative improvements on solution accuracy and efficiency from SOTA normal- and behavioral-form solvers (CFR+, MMD, Deep-CFR, JPSPG, PPO, R-NaD). (2) We further demonstrate the practical value of the equilibrium structure of 2p01s by solving an American football setting where the attacking team needs to strategically conceal their true intention between “RB power push” and “QB throw”. While this IIEFG has a complexity of $10^{440}$ even with a coarse mesh of 10 values per action dimension, the atomic structure of its NE makes the game solvable in less than 30 minutes on a M1 Pro Macbook. See Fig. 1(b,c).
@@ 2 Related Work
Games where players have missing information only about the game types are called incomplete-information. These are a special case of imperfect-information games where nature only plays a chance move at the beginning Harsanyi (1967). The seminal work of Aumann et al. (1995) developed equilibrium strategies for a repeated game with one-sided incomplete information through the “Cav u” theorem, which reveals that belief-manipulating behavioral strategies optimize value through value convexification. Building on top of this, De Meyer (1996) introduced a dual game reformulation where the behavioral strategy of P2 becomes Markov. This technique helped Cardaliaguet (2007); Ghimire et al. (2024) to establish the value existence proof for 2p0s differential games with incomplete information. Unlike repeated games where belief manipulation occurs only in the first round of the game, differential games may have multiple critical time-state-belief points where belief manipulation is required to achieve equilibrium, depending on the specifications of system dynamics, payoffs, and state constraints (Ghimire et al., 2024). Our paper builds on top of Cardaliaguet (2007); Ghimire et al. (2024), providing a thorough analysis of the convexification nature of behavioral strategies and, as a consequence, their atomic structure.
@@
IIEFGs are partially-observable stochastic games (POSGs) with finite horizons. Significant efforts have been taken to approximate equilibrium of large IIEFGs with finite $U$ (Koller & Megiddo, 1992; Billings et al., 2003; Gilpin & Sandholm, 2006; Gilpin et al., 2007; Sandholm, 2010; Brown & Sandholm, 2019; Brown et al., 2020a; Perolat et al., 2022; Schmid et al., 2023). Regret matching algorithms (Zinkevich et al., 2007; Lanctot et al., 2009; Abernethy et al., 2011; Brown et al., 2019; Tammelin, 2014; Johanson et al., 2012) have computational complexity $\mathcal{O}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U}\varepsilon^{-2})$ to $\varepsilon$-Nash, while gradient-based solvers (McMahan, 2011; Perolat et al., 2021; Sokota et al., 2022) have $\mathcal{O}\left(\ln({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U})\varepsilon^{-1}\ln\left(\varepsilon^{-1}\right)\right)$ to $\varepsilon$-QRE. All these complexities increase with $U$, and convergence guarantee only exists if the NE lies in the interior of the simplex $\Delta(U)$. Critically, this latter assumption does not hold for 2p0s1, as we explain in Sec. 4. For games with continuous action spaces, existing pseudo-gradient based approaches (Martin & Sandholm, 2023, 2024) lack convergence guarantee, and perform poorly in our case studies in both normal- and extensive-form settings.
@@
Since solving 2p0s1 essentially requires solving a Hamilton-Jacobi PDE, we briefly review multigrid methods for accelerating PDE solving(Trottenberg et al., 2000). In a typical V-cycle solver (Braess & Hackbusch, 1983), a fine-mesh solve is first performed briefly, the residual is then restricted to a coarser mesh where a correction is solved and prolongated back to the fine mesh. Essentially, multigrid uses coarse solves to reduce the low-frequency approximation errors in the PDE solution at low costs, leaving only high-frequency errors to be resolved through the fine mesh. Multigrid was successfully applied to solving Hamilton-Jacobi PDEs for optimal control and differential games (Han & Wan, 2013), although characterizing its computational complexity for such nonlinear PDEs is yet to be done. During value approximation, we extend multigrid to solving Hamilton-Jacobi PDEs underlying 2p0s1.
@@ 3 Problem Statement
The initial belief of players is set to nature’s distribution about the game type. Denote by $\{\mathcal{H}^{i}_{r}\}^{I}$ the joint sets of $I$ behavioral strategies of P1, and $\mathcal{Z}_{r}$ the set of behavioral strategies of P2. The subscript $r$ indicates the random nature of behavioral strategies: At any infostate $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$ and with type $i$, P1 draws an action based on $\eta_{i}\in\mathcal{H}_{r}^{i}$, which is a probability measure over $\mathcal{U}$ parameterized by $(t,x,p)$. P2’s strategy $\zeta\in\mathcal{Z}_{r}(t)$ is a probability measure over $\mathcal{V}$ independent of $i$. Denote by $X_{t_{1}}^{t_{0},x_{0},\eta_{i},\zeta}$ the random state arrived at $t_{1}\in(t_{0},T]$ from $(t_{0},x_{0})$ following $(\eta_{i},\zeta)$ and the system dynamics in Eq. 1. With mild abuse of notation, let $(\eta_{i}(t),\zeta(t))$ denote the random controls at time $t$ induced by $(\eta_{i},\zeta)$. P1 accumulates a running cost $l_{i}(\eta_{i}(t),\zeta(t))$ during the game and receives a terminal cost $g_{i}(X_{T}^{t_{0},x_{0},\eta_{i},\zeta})$. Together, a type-$i$ P1 minimizes:
$J_{i}(t_{0},x_{0};\eta_{i},\zeta):=\mathbb{E}_{\eta_{i},\zeta}\left[g_{i}\left(X_{T}^{t_{0},x_{0},\eta_{i},\zeta}\right)+\int_{t_{0}}^{T}l_{i}(\eta_{i}(s),\zeta(s))ds\right],$
while P2 maximizes $J(t_{0},x_{0},p_{0};\{\eta_{i}\},\zeta)=\mathbb{E}_{i\sim p_{0}}[J_{i}]$. We say the game has a value $V$ if and only if the upper value $V^{+}(t_{0},x_{0},p_{0})=\inf_{\{\eta_{i}\}}\sup_{\zeta}J$ and the lower value $V^{-}(t_{0},x_{0},p_{0})=\sup_{\zeta}\inf_{\{\eta_{i}\}}J$ are equal: $V=V^{+}=V^{-}$. The game is proven to have a value under the following sufficient conditions (Cardaliaguet, 2007):
@@
    $g_{i}:\mathcal{X}\rightarrow\mathbb{R}$ and $l_{i}:\mathcal{U}\times\mathcal{V}\rightarrow\mathbb{R}$ are bounded and Lipschitz continuous.
@@
    Isaacs’ condition holds for the Hamiltonian $H:\mathcal{X}\times\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}$:
    $H(x,\xi)$ $=\min_{u\in\mathcal{U}}\max_{v\in\mathcal{V}}f(x,u,v)^{\top}\xi+l_{i}(u,v)=\max_{v\in\mathcal{V}}\min_{u\in\mathcal{U}}f(x,u,v)^{\top}\xi+l_{i}(u,v).$ (2)
@@
Our goal is to compute a Nash equilibrium (NE) $(\{\eta_{i}\}^{\dagger},\zeta^{\dagger})$ that attains $V$, given the game $G=\{\mathcal{X},(\mathcal{U},\mathcal{V}),(\{g_{i}\},\{l_{i}\}),f,T\}$.
@@ 4 A Primal-Dual Reformulation of the Game
$G_{\tau}$ is a discrete-time Stackelberg version of $G$ where P2 plays a pure best response after P1 announces his next action. Let the value of $G_{\tau}$ be $V_{\tau}$. $V_{\tau}$ satisfies the following DP for $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$:
$V_{\tau}(t,x,p)=\min_{\{\eta_{i}\}\in\{\mathcal{H}_{r}\}^{I}}\mathbb{E}_{i\sim p,u\sim\eta_{i}}\Big[$ $\max_{v\in\mathcal{V}}V_{\tau}(t+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau l_{i}(u,v)\Big],$ (3)
with a terminal boundary $V_{\tau}(T,x,p)=\sum_{i}p[i]g_{i}(x)$. For small enough $\tau$, $x^{\prime}(u,v)=x+\tau f(x,u,v)$. $p^{\prime}(u)$ is the Bayes update of the public belief after P1 announces $u$: $p^{\prime}(u)[i]=\eta_{i}(u;t,x,p)p[i]/\bar{\eta}(u;t,x,p)$, where $\bar{\eta}(u;t,x,p)=\sum_{i\in[I]}\eta_{i}(u;t,x,p)p_{0}[i]$ is the marginal over $\mathcal{U}$ across types. Note that P2’s equilibrium behavioral strategy cannot be derived from Eq. 3.
@@
For P2’s strategy, we need a separate DP where P2 announces his next action and P1 best responses to it. We do so by first introducing the convex conjugate $V^{*}$ of the value:
@@
Eq. 4 describes a dual game $G^{*}$ with complete information, where the strategy space of P1 becomes $\mathcal{H}\times[I]$, i.e., the game type is now chosen by P1 rather than the nature. We prove in App. A that P2’s equilibrium in the dual game is also an equilibrium in the primal game if $\hat{p}_{0}\in\partial_{p}V(t_{0},x_{0},p_{0})$, and such $\hat{p}_{0}[i]$ represents the loss of type-$i$ P1 should he play best responses to P2’s equilibrium strategy. Therefore $\hat{p}_{0}[i]-\mathbb{E}_{\zeta}[J_{i}]$ measures P2’s risk, and P2’s equilibrium strategy minimizes the worst-case risk across all game types. We now introduce a discrete-time version of the dual game $G^{*}_{\tau}$ where P1 plays a pure best response after P2 announces their action at each time step. Let the value of $G^{*}_{\tau}$ be $V_{\tau}^{*}$, the dual DP is:
$V_{\tau}^{*}(t,x,\hat{p})=\min_{\zeta,\hat{p}^{\prime}(\cdot)}\mathbb{E}_{v\sim\zeta}\left[\max_{u\in\mathcal{U}}V_{\tau}^{*}(t+\tau,x^{\prime}(u,v),\hat{p}^{\prime}(v)-\tau l(u,v))\right],$ (5)
@@
$\min_{\{u^{k}\},\{\alpha_{i}^{k}\}}\max_{\{v^{k}\}}\quad\sum_{k=1}^{I}\lambda^{k}\Big(V(t+\tau,x+\tau f(x,u^{k},v^{k}),p^{k})+\tau\mathbb{E}_{i\sim p^{k}}[l_{i}(u^{k},v^{k})]\Big)$ ($\text{P}_{1}$)
@@
$\min_{\{v^{k}\},\{\lambda^{k}\},\{\hat{p}^{k}\}}\max_{\{u^{k}\}}\sum_{k=1}^{I+1}\lambda^{k}\left(V^{*}(t+\tau,x+\tau f(x,u^{k},v^{k}),\hat{p}^{k}-\tau l(u^{k},v^{k}))\right)$ ($\text{P}_{2}$)
@@
To support the proof sketch, we use Fig. 2 to illustrate why an equilibrium behavioral strategy is atomic. Here $I=2$, thus the value is defined on $\Delta(2)$ for fixed $(t,x)$. To solve $\min_{\nu}\int_{\Delta(2)}\tilde{V}_{\tau}(p^{\prime})\nu(dp^{\prime})$, we scan the non-revealing value $\tilde{V}$ across $\Delta(2)$. One notices that if $\tilde{V}$ is not convex in $p$, it is always possible for P1 to achieve a lower loss by convexifying $\tilde{V}$ through a mixed strategy, leading to $\text{P}_{1}$. In the figure, P1 identifies $\lambda^{a},\lambda^{b}]^{\top}\in\Delta(2)$ and $\{p^{a},p^{b}\}$ such that $\lambda^{a}p^{a}+\lambda^{b}p^{b}=p$. Solving the non-revealing pure NE $(u^{k},v^{k})$ for each $k\in\{a,b\}$, P1’s mixed strategy that convexifies $\tilde{V}$ is to play action $u^{k}$ with probability $\alpha^{k}_{i}=p^{k}[i]\lambda^{k}/p[i]$ if he is type-$i$. By announcing this strategy, the public belief shifts to $p^{k}$ via the Bayes’ rule when P2 observes action $u^{k}$. As a result, P1 receives $V(p)=\lambda^{a}\tilde{V}(p^{a})+\lambda^{b}\tilde{V}(p^{b})$ on the convex hull of $\tilde{V}(p)$.
@@
(1) The atomic structure of the equilibrium strategies was first discovered for 2p0s repeated games with one-sided information (Aumann et al., 1995; De Meyer, 1996). Our new contribution is in explaining its presence in differential games and in demonstrating its significant utilities in improving the effectiveness of multiagent reinforcement learning and model predictive control schemes. (2) P1’s actions in 2p0s1 simultaneously advance system states and achieve signaling. The deterministic dynamics allows precise belief control (e.g., $(p^{a},p^{b})$ in Fig. 2) for value convexification. For games with stochastic dynamics/observations, however, P1 will not be able to precisely control the belief, and the convex Bellman operators ($\text{P}_{1}$ and $\text{P}_{2}$) become lower bounds of the value. Finding tight upper-bounding operators that preserve atomic NEs is left for future work.

### v4

@@ Abstract
For a two-player imperfect-information extensive-form game (IIEFG) with $K$ time steps and a player action space of size $U$, the game tree complexity is $U^{2K}$, causing existing IIEFG solvers to struggle with large or infinite $(U,K)$, e.g., differential games with continuous action spaces. To partially address this scalability challenge, we focus on an important class of 2p0s games where the informed player (P1) knows the payoff while the uninformed player (P2) only has a belief over the set of $I$ possible payoffs. Such games encompass a wide range of scenarios in sports, defense, cybersecurity, and finance. We prove that under mild conditions, P1’s (resp. P2’s) equilibrium strategy at any infostate concentrates on at most $I$ (resp. $I+1$) action prototypes. When $I\ll U$, this equilibrium structure causes the game tree complexity to collapse to $I^{K}$ for P1 when P2 plays best responses, and $(I+1)^{K}$ for P2 in a dual game where P1 plays best responses. We then show that exploiting this structure in model-free multiagent reinforcement learning and model predictive control leads to significant improvements in learning accuracy and efficiency from SOTA IIEFG solvers. Our demonstration solves a 22-player football game with continuous action spaces and $K=10$ time steps, where the offense team needs to strategically conceal their play until a critical moment in order to exploit information advantage. Code is available here.
@@ 1 Introduction
The strength of game solvers has grown rapidly in the last decade, beating elite-level human players in Chess (alphazero), Go (alphago), Poker (pluribus; rebel), Diplomacy (diplomacy), Stratego (stratego), among others with increasing complexity. However, most of the existing solvers with proven convergence, either based on regret matching (tammelin2014solving; burch2014solving; moravvcik2017deepstack; rebel; lanctot2009monte) or gradient descent-ascent (pmlr-v15-mcmahan11b; perolat2021poincare; sokota2022unified; cen2021fast; vieillard2020leverage), have computational complexities increasing with the size of the finite action set, and suffer from game-tree complexities growing exponentially with both the action size $U$ and the tree depth $K$. Real-world games, however, often have continuous actions and happen in continuous time and state spaces, making them differential in nature. Applying existing solvers to differential games would require game-specific insight or extra computational overhead for automated abstraction (kroer2015discretization; hawkin2011automated; brown2015simultaneous).
@@
In this paper, we address this scalability challenge for an K subset of 2p0s differential games where the informed player (P1) knows the payoff type while the uninformed player (P2) only holds a public belief $p_{0}\in\Delta(I)$ over the finite set of $I$ possible types. At the beginning of the game, nature draws a game type according to $p_{0}$ and informs P1 about the type. As the game progresses, the public belief $p$ about the true game type is updated from $p_{0}$ based on the action sequence taken by P1 and its strategy via the Bayes’ rule. P1’s (resp. P2’s) goal is to minimize (resp. maximize) the expected payoff over $p_{0}$. Due to the zero-sum nature, P1 may need to delay information release or manipulate P2’s belief to take full advantage of information asymmetry. While restricted, such games represent a wide range of attack-defense scenarios including football set-pieces where the attacker has private information about which play is to be executed, and missile defense where multiple potential targets are concerned. The setting of one-sided information, i.e., P1 knows everything about P2, is necessary for P2 to derive defense strategies in risk-sensitive games. We call this focused set of games “2p0s1”.
@@
  • We demonstrate how this structural knowledge can significantly accelerate game solving: (1) For value and policy approximation settings where the ground-truth NE is available, we achieve qualitative improvements on solution accuracy and efficiency from SOTA normal- and behavioral-form solvers (CFR+, MMD, Deep-CFR, JPSPG, PPO, R-NaD). (2) We further demonstrate the practical value of the equilibrium structure of 2p0s1 by solving an American football setting where the attacking team needs to strategically conceal their true intention between “RB power push” and “QB throw”. While this IIEFG has a complexity of $10^{440}$ even with a coarse mesh of 10 values per action dimension, the atomic structure of its NE makes the game solvable in less than 30 minutes on a M1 Pro Macbook. See Fig. 1b,c.
@@ 2 Related Work
Games where players have missing information only about the game types are called incomplete-information. These are a special case of imperfect-information games where nature only plays a chance move at the beginning harsanyi1967games. The seminal work of aumann1995repeated developed equilibrium strategies for a repeated game with one-sided incomplete information through the “Cav u” theorem, which reveals that belief-manipulating behavioral strategies optimize value through value convexification. Building on top of this, de1996repeated introduced a dual game reformulation where the behavioral strategy of P2 becomes Markov. This technique helped cardaliaguet2007differential; ghimire24a to establish the value existence proof for 2p0s differential games with incomplete information. Unlike repeated games where belief manipulation occurs only in the first round of the game, differential games may have multiple critical time-state-belief points where belief manipulation is required to achieve equilibrium, depending on the specifications of system dynamics, payoffs, and state constraints (ghimire24a). Our paper builds on top of cardaliaguet2007differential; ghimire24a, providing a thorough analysis of the convexification nature of behavioral strategies and, as a consequence, their atomic structure.
@@
IIEFGs are partially-observable stochastic games (POSGs) with finite horizons. Significant efforts have been taken to approximate equilibrium of large IIEFGs with finite $U$ (koller1992complexity; billings2003approximating; gilpin2006finding; gilpin2007gradient; sandholm2010state; pluribus; brown2020combining; stratego; sog). Regret matching algorithms (zinkevich2007regret; lanctot2009monte; abernethy2011blackwell; brown2019deep; tammelin2014solving; johanson2012finding) have computational complexity $\mathcal{O}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U}\varepsilon^{-2})$ to $\varepsilon$-Nash, while gradient-based solvers (pmlr-v15-mcmahan11b; perolat2021poincare; sokota2022unified) have $\mathcal{O}\left(\ln({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U})\varepsilon^{-1}\ln\left(\varepsilon^{-1}\right)\right)$ to $\varepsilon$-QRE. All these complexities increase with $U$, and convergence guarantee only exists if the NE lies in the interior of the simplex $\Delta(U)$ (perolat2021poincare). Critically, this latter assumption does not hold for 2p0s1, as we explain in Sec. 4. For games with continuous action spaces, existing pseudo-gradient based approaches (martin2023finding; martin2024joint) lack convergence guarantee, and perform poorly in our case studies in both normal- and extensive-form settings.
@@
Since solving 2p0s1 essentially requires solving a Hamilton-Jacobi PDE, we briefly review multigrid methods for accelerating PDE solving(trottenberg2000multigrid). In a typical V-cycle solver (braess1983new), a fine-mesh solve is first performed briefly, the residual is then restricted to a coarser mesh where a correction is solved and prolongated back to the fine mesh. Essentially, multigrid uses coarse solves to reduce the low-frequency approximation errors in the PDE solution at low costs, leaving only high-frequency errors to be resolved through the fine mesh. Multigrid was successfully applied to solving Hamilton-Jacobi PDEs for optimal control and differential games (han2013multigrid), although characterizing its computational complexity for such nonlinear PDEs is yet to be done. During value approximation, we extend multigrid to solving Hamilton-Jacobi PDEs underlying 2p0s1.
@@ 3 Problem Statement
The initial belief of players is set to nature’s distribution about the game type. Denote by $\{\mathcal{H}^{i}_{r}\}^{I}$ the joint sets of $I$ behavioral strategies of P1, and $\mathcal{Z}_{r}$ the set of behavioral strategies of P2. The subscript $r$ indicates the random nature of behavioral strategies: At any infostate $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$ and with type $i$, P1 draws an action based on $\eta_{i}\in\mathcal{H}_{r}^{i}$, which is a probability measure over $\mathcal{U}$ parameterized by $(t,x,p)$. P2’s strategy $\zeta\in\mathcal{Z}_{r}$ is a probability measure over $\mathcal{V}$ independent of $i$. Denote by $X_{t_{1}}^{t_{0},x_{0},\eta_{i},\zeta}$ the random state arrived at $t_{1}\in(t_{0},T]$ from $(t_{0},x_{0})$ following $(\eta_{i},\zeta)$ and the system dynamics in Eq. 1. With mild abuse of notation, let $(\eta_{i}(t),\zeta(t))$ denote the random controls at time $t$ induced by $(\eta_{i},\zeta)$. P1 accumulates a running cost $l_{i}(x(t),\eta_{i}(t),\zeta(t))$ during the game and receives a terminal cost $g_{i}(X_{T}^{t_{0},x_{0},\eta_{i},\zeta})$. Together, a type-$i$ P1 minimizes:
$J_{i}(t_{0},x_{0};\eta_{i},\zeta):=\mathbb{E}_{\eta_{i},\zeta}\left[g_{i}\left(X_{T}^{t_{0},x_{0},\eta_{i},\zeta}\right)+\int_{t_{0}}^{T}l_{i}(x(s),\eta_{i}(s),\zeta(s))ds\right],$
while P2 maximizes $J(t_{0},x_{0},p_{0};\{\eta_{i}\},\zeta)=\mathbb{E}_{i\sim p_{0}}[J_{i}]$. We say the game has a value $V$ if and only if the upper value $V^{+}(t_{0},x_{0},p_{0})=\inf_{\{\eta_{i}\}}\sup_{\zeta}J$ and the lower value $V^{-}(t_{0},x_{0},p_{0})=\sup_{\zeta}\inf_{\{\eta_{i}\}}J$ are equal: $V=V^{+}=V^{-}$. The game is proven to have a value under the following sufficient conditions (cardaliaguet2007differential):
@@
    $g_{i}:\mathcal{X}\rightarrow\mathbb{R}$ and $l_{i}:\mathcal{X}\times\mathcal{U}\times\mathcal{V}\rightarrow\mathbb{R}$ are bounded and Lipschitz continuous.
@@
    Isaacs’ condition holds for the Hamiltonian $H:\mathcal{X}\times\mathbb{R}^{d_{x}}\times\Delta(I)\rightarrow\mathbb{R}$:
    $H(x,\xi,p)$ $=\min_{u\in\mathcal{U}}\max_{v\in\mathcal{V}}f(x,u,v)^{\top}\xi+\mathbb{E}_{i\sim p}[l_{i}(x,u,v)]$ (2)
    $=\max_{v\in\mathcal{V}}\min_{u\in\mathcal{U}}f(x,u,v)^{\top}\xi+\mathbb{E}_{i\sim p}[l_{i}(x,u,v)].$
@@
Our goal is to compute a Nash equilibrium (NE) $(\{\eta_{i}^{\dagger}\},\zeta^{\dagger})$ that attains $V$, given the game $G=\{\mathcal{X},(\mathcal{U},\mathcal{V}),\{g_{i}\},\{l_{i}\},f,T\}$.
@@ 4 A Primal-Dual Reformulation of the Game
$G_{\tau}$ is a discrete-time Stackelberg version of $G$ where P2 plays a pure best response after P1 announces its next action. Let the value of $G_{\tau}$ be $V_{\tau}$. $V_{\tau}$ satisfies the following DP for $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$:
$V_{\tau}(t,x,p)=\min_{\{\eta_{i}\}\in\{\mathcal{H}_{r}\}^{I}}\mathbb{E}_{i\sim p,u\sim\eta_{i}}\Big[$ $\max_{v\in\mathcal{V}}V_{\tau}(t+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau l_{i}(x,u,v)\Big],$ (3)
with a terminal boundary $V_{\tau}(T,x,p)=\sum_{i}p[i]g_{i}(x)$. For small enough $\tau$, $x^{\prime}(u,v)=x+\tau f(x,u,v)$. $p^{\prime}(u)$ is the Bayes update of the public belief after P1 announces $u$: $p^{\prime}(u)[i]=\eta_{i}(u;t,x,p)p[i]/\bar{\eta}(u;t,x,p)$, where $\bar{\eta}(u;t,x,p)=\sum_{i\in[I]}\eta_{i}(u;t,x,p)p[i]$ is the marginal over $\mathcal{U}$ across types. Note that P2’s equilibrium behavioral strategy cannot be derived from Eq. 3.
@@
For P2’s strategy, we need a separate DP where P2 announces its next action and P1 best responses to it. We do so by first introducing the convex conjugate $V^{*}$ of the value:
@@
Eq. 4 describes a dual game $G^{*}$ with complete information, where the strategy space of P1 becomes $\mathcal{H}\times[I]$, i.e., the game type is now chosen by P1 rather than the nature. We prove in App. A that P2’s equilibrium in the dual game is also an equilibrium in the primal game if $\hat{p}_{0}\in\partial_{p}V(t_{0},x_{0},p_{0})$, and such $\hat{p}_{0}[i]$ represents the loss of type-$i$ P1 should it play best responses to P2’s equilibrium strategy. Therefore $\hat{p}_{0}[i]-\mathbb{E}_{\zeta}[J_{i}]$ measures P2’s risk, and P2’s equilibrium strategy minimizes the worst-case risk across all game types. We now introduce a discrete-time version of the dual game $G^{*}_{\tau}$ where P1 plays a pure best response after P2 announces their action at each time step. Let the value of $G^{*}_{\tau}$ be $V_{\tau}^{*}$, the dual DP is:
$V_{\tau}^{*}(t,x,\hat{p})=\min_{\zeta,\hat{p}^{\prime}(\cdot)}\mathbb{E}_{v\sim\zeta}\left[\max_{u\in\mathcal{U}}V_{\tau}^{*}(t+\tau,x^{\prime}(u,v),\hat{p}^{\prime}(v)-\tau l(x,u,v))\right],$ (5)
@@
$\min_{\{u^{k}\},\{\alpha_{i}^{k}\}}\max_{\{v^{k}\}}\quad\sum_{k=1}^{I}\lambda^{k}\Big(V(t+\tau,x+\tau f(x,u^{k},v^{k}),p^{k})+\tau\mathbb{E}_{i\sim p^{k}}[l_{i}(x,u^{k},v^{k})]\Big)$ ($\text{P}_{1}$)
@@
$\min_{\{v^{k}\},\{\lambda^{k}\},\{\hat{p}^{k}\}}\max_{\{u^{k}\}}\sum_{k=1}^{I+1}\lambda^{k}\left(V^{*}(t+\tau,x+\tau f(x,u^{k},v^{k}),\hat{p}^{k}-\tau l(x,u^{k},v^{k}))\right)$ ($\text{P}_{2}$)
@@
To support the proof sketch, we use Fig. 2 to illustrate why an equilibrium behavioral strategy is atomic. Here $I=2$, thus the value is defined on $\Delta(2)$ for fixed $(t,x)$. To solve $\min_{\nu}\int_{\Delta(2)}\tilde{V}_{\tau}(p^{\prime})\nu(dp^{\prime})$, we scan the non-revealing value $\tilde{V}$ across $\Delta(2)$. One notices that if $\tilde{V}$ is not convex in $p$, it is always possible for P1 to achieve a lower loss by convexifying $\tilde{V}$ through a mixed strategy, leading to $\text{P}_{1}$. In the figure, P1 identifies $\lambda^{a},\lambda^{b}]^{\top}\in\Delta(2)$ and $\{p^{a},p^{b}\}$ such that $\lambda^{a}p^{a}+\lambda^{b}p^{b}=p$. Solving the non-revealing pure NE $(u^{k},v^{k})$ for each $k\in\{a,b\}$, P1’s mixed strategy that convexifies $\tilde{V}$ is to play action $u^{k}$ with probability $\alpha^{k}_{i}=p^{k}[i]\lambda^{k}/p[i]$ if P1 is type-$i$. By announcing this strategy, the public belief shifts to $p^{k}$ via the Bayes’ rule when P2 observes action $u^{k}$. As a result, P1 receives $V(p)=\lambda^{a}\tilde{V}(p^{a})+\lambda^{b}\tilde{V}(p^{b})$ on the convex hull of $\tilde{V}(p)$.
@@
(1) The atomic structure of the equilibrium strategies was first discovered for 2p0s repeated games with one-sided information (aumann1995repeated; de1996repeated). Our new contribution is in explaining its presence in differential games and in demonstrating its significant utilities in improving the effectiveness of multiagent reinforcement learning and model predictive control schemes. (2) P1’s actions in 2p0s1 simultaneously advance system states and achieve signaling. The deterministic dynamics allows precise belief control (e.g., $(p^{a},p^{b})$ in Fig. 2) for value convexification. For games with stochastic dynamics/observations, however, P1 will not be able to precisely control the belief, and the convex Bellman operators ($\text{P}_{1}$ and $\text{P}_{2}$) become lower bounds of the value. Finding tight upper-bounding operators that preserve atomic NEs is left for future work.